OpenStudy (anonymous):

let sinx= 15/17, cosx<0, tany=-12/5 andy siny<0, find sin(x+y)

OpenStudy (anonymous):

i know that sin(x+y) = sinxcosy+cosxsiny =15/17cosy + cosxsiny

OpenStudy (anonymous):

15/17cosy + cosxsiny=cos y (15/17+cos x tan y)=cos y (15/17-12/5cos x) so u just need to evaluate cos y and cos x

OpenStudy (anonymous):



OpenStudy (anonymous):

Find cos x = 8/17

OpenStudy (anonymous):

use this identities \[\sin^2 x+\cos^2 x=1\]\[1+\tan^2 y=\frac{1}{\cos^2 y}\]

OpenStudy (anonymous):

draw your triangles... for angle y, you have a negative tangent so so cosy will be postive....

OpenStudy (anonymous):

cos x = -8/17 because cos x < 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do the same thing that @Yahoo! did with the other angle, angle y.....

OpenStudy (anonymous):